Let S = $52, s = 20%, and r = 7% (continuously compounded). The stock is set to pay a single dividend of $1.10 nine months from today, with no further dividends expected this year. Use the Black-Scholes model (adjusted for the dividend) to compute the value of a one-year $50-strike European call option on the stock.

answer= $6.43

Please show all the steps. Thanks
Answer #1

Interest Rate = r = 7% or 0.07/12 monthly

Dividends received in none months = D9 = 1.10

Present Value of Dividend = D0 = D9/(1+r)n = 1.10/(1+0.07/12)9 = 1.044

S = Dividend adjusted Stock Price = 52 - 1.044 50.956
t = time until option expiration(years) = 1
K = Option Strike Price = 50
r = risk free rate(annual) = 7% = 0.07
s = standard deviation(annual) = 0.20
N = cumulative standard normal distribution
d1 = {ln (S/K) + (r +s^2/2)t}/s√t
= {ln (50.956/50) + (0.07 + 0.2^2/2)*1}/0.2*√1
0.5447
d2 = d1 - s√t
= 0.5447 - 0.2√1
0.3447
Using z tables,
N(d1) = 0.7070
N(d2) = 0.6348
C = Call Premium = =SN(d1) - N(d2)Ke^(-rt)
= 50.956*0.707 - 0.6348*50e^(-0.07*1)
6.4317

Hence, Value of call option = C = $6.43